# PerLA  [Uncaptioned image] : Perceptive 3D language assistant

### v1

@@ Abstract
Enabling Large Language Models (LLMs) to understand the 3D physical world is an emerging yet challenging research direction. Current strategies for processing point clouds typically downsample the scene or divide it into smaller parts for separate analysis. However, both approaches risk losing key local details or global contextual information. In this paper, we introduce PerLA, a 3D language assistant designed to be more perceptive to both details and context, making visual representations more informative for the LLM. PerLA captures high-resolution (local) details in parallel from different point cloud areas and integrates them with (global) context obtained from a lower-resolution whole point cloud. We present a novel algorithm that preserves point cloud locality through the Hilbert curve and effectively aggregates local-to-global information via cross-attention and a graph neural network. Lastly, we introduce a novel loss for local representation consensus to promote training stability. PerLA outperforms state-of-the-art 3D language assistants, with gains of up to +1.34 CiDEr on ScanQA for question answering, and +4.22 on ScanRefer and +3.88 on Nr3D for dense captioning. [^1] Project page: https://gfmei.github.io/PerLA/
Corresponding author: gmei@fbk.eu
@@ 2 Related work
3D language understanding involves understanding 3D scenes by describing or answering to scene-relevant questions in the format of natural language. Examples of typical downstream tasks include 3D Dense Captioning [8, 12, 53], and 3D Question Answering [2, 36, 61, 63]. 3D Dense Captioning consists of 3D localization of object instances, and textual descriptions of each object instance [12, 4, 8, 10]. 3D Question Answering requires the model equipped with a language decoder to answer questions regarding the visual context in the given 3D scene [2, 41]. Several works focus on addressing the problems of 3D-language pre-alignment [13, 25], or designing adapter layers [21, 9], or building 3D sintetic data [60]. In contrast, our work focuses on designing an encoding approach that can enhance 3D-language understanding by capturing fine-grained details.
3D language assistants (3DLAs) can be roughly categorized into two types: object-level 3DLAs [43, 50, 57] and scene-level 3DLAs [21, 65, 55, 9]. Object-level 3DLAs learn from large 3D object datasets [15] to connect object-level 3D representation with language models. However, they underperform in compositional reasoning in complex 3D scenes with numerous objects. Scene-level 3DLAs, such as 3D-LLM [21], LLaVA-3D [65], Chat-3D [55], and LL3DA [9] enable scene understanding via the interaction with objects. Because 3D data is orders of magnitude less than 2D data, existing 3DLAs address such lack of data by leveraging pre-trained 2D Large Multimodal Models (LMMs) [21, 65], or data-efficient training recipes [55, 9]. 3D-LLM [21] leverages 2D pre-trained representations of rendered multi-view images to construct 3D representations and 2D VLMs as backbones. LLaVA-3D [65] adapts LLaVA [33] for 3D scene understanding by associating 2D patch representations with their positions in 3D space. Chat-Scene [22] improves the referencing and grounding capability and models scene representations as a sequence of object-level representations. LL3DA [9] extracts point-level representations from a downsampled 3D scene, and includes both interaction prompts and textual instructions to resamble human interactions with the 3D environment. Orthogonal to the efforts of aligning 3D-language with limited 3D data in existing 3DLAs, PerLA aims to improve the capability of 3DLAs in perceiving scene details.
@@ 3.1 Perceptive scene encoder
Global and partial scene encoding. Both full-scene point cloud and the partial-scene point clouds are encoded separately by the same pre-trained 3D scene encoder $\phi$. The 3D scene encoder $\phi$ downsamples the input point cloud into super-points using Farthest Point Sampling (FPS) [38] and produces a representation for each super-point. The representation of the full scene, i.e.the global representation, encode the overall scene context, while the representations of the partial scenes, i.e.the local representations, encode scene details. Let $\mathcal{P}^{g}=\{\bm{p}^{g}_{i}\in\mathbb{R}^{3}\}_{i=1}^{M}$ denote the downsampled full-scene point cloud composed of $M$ points, and $\mathcal{F}^{g}=\{\bm{f}^{g}_{i}\in\mathbb{R}^{d}\}_{i=1}^{M}$ the associated global representations. Let $\mathcal{P}^{l}=\{\bm{p}^{l}_{i}\in\mathbb{R}^{3}\}_{i=1}^{L\cdot M}$ denote the set of downsampled partial-scene point clouds, and $\mathcal{F}^{l}=\{\bm{f}^{l}_{i}\in\mathbb{R}^{d}\}^{L\cdot M}$ the associated local representations. Note that, we aggregate all the points of the downsampled partial-scene point clouds into the single set $\mathcal{P}^{l}$, and both the full-scene and the partial-scenes are downsampled to the same number of super-points $M$. The representations $\mathcal{F}^{l}$ provide a more detailed view of the scene, as they are derived from the same number of points downsampled from smaller, localized regions, resulting in higher resolution compared to the global versions. Yet, their semantic visibility is focused within each local part. In the following steps, we aggregate local and global information, to enrich the representations with details within the scene context.
@@
Hilbert-based nearest-neighbor search. In order to enhance global representations through local representations, we need to first find the correspondences among them, i.e.which super-points of partial-scene point clouds are spatially neighbors with super-points of the full-scene point cloud. This can be done via $k$-NN search. However, traditional $k$-NN searches can be computationally intensive on large-scale or high-resolution point clouds. To address this, we approximate $k$-NN with an efficient neighbor-mapping technique using Hilbert-serialized point clouds, which enhances query speed while reducing computational complexity. With serialized point clouds, we can leverage point indices to perform $k$-NN searches with $O(1)$ complexity [56].
@@
Specifically, we first serialize the union $\mathcal{P}^{g}\bigcup\mathcal{P}^{l}$ using Hilbert curve ordering. Then, for each $\bm{p}^{g}_{i}\in\mathcal{P}^{g}$, we identify the k nearest super-points within $\mathcal{P}^{l}$ based on the serialized indexes. Let $\mathcal{P}_{\texttt{k}_{i}}^{l}$ denote the set of k nearest local super-points to the global super-point $\bm{p}^{g}_{i}$. Consequently, $\mathcal{F}_{\texttt{k}_{i}}^{l}$ are the representations of the super-points in $\mathcal{P}_{\texttt{k}_{i}}^{l}$.
@@
Localized cross-attention operates with a learning-based weighting mechanism that is a function of the representations and relative positions of global super-points and their associated local neighbors. We constrain cross-attention to local neighborhood regions because i) limiting the number of points reduces computational complexity, and ii) neighbors of a given point are likely to belong to the same object. Specifically, for each point $\bm{p}^{g}_{i}$ and its nearest-neighbors $\mathcal{P}_{\texttt{k}_{i}}^{l}$, their relative position embedding $\mathcal{R}_{i}$ are extracted as
$\vspace{-2mm}\mathcal{R}_{i}=\left\{\mathcal{R}_{ij}=\text{pos}\left(\left(\bm{p}^{g}_{i}-\bm{p}^{l}_{j}\right)/{\sigma}\right)\in\mathbb{R}^{d}\right\}_{j=1}^{K^{l}},$ (1)
@@
$\text{pos}(\bm{x})=\left[\sin\left(2\pi\bm{x}\cdot B\right);\cos\left(2\pi\bm{x}\cdot B\right)\right],$ (2)
@@
GCN message passing further refines ^$\hat{\mathcal{F}}^{g}$ by aggregating information from neighboring global super-points. Let $\mathcal{G}^{g}=\{\mathcal{P}^{g},\mathcal{W}^{g}\}$ be a $k$-NN graph, where $\mathcal{P}^{g}$ is the set of the vertices, i.e. global super-points, and $\mathcal{W}^{g}\in\mathbb{R}^{N\times N}$ is the adjacency matrix defining the edges. The construction of the adjacency matrix uses the similarity between super-point representations $\mathcal{F}^{G}$. Note that we use the (original) global representations to define the adjacency matrix because we empirically experienced more stability in training as opposed to using the updated global representation, likely due to the fact that they change value during training. The adjacency matrix $\mathcal{W}^{g}$ with elements $\mathcal{W}^{g}_{ij}$ is defined as
$\mathcal{W}^{g}_{ij}{=}\begin{cases}\frac{{\bm{f}^{g}_{i}}^{\top}\bm{f}^{g}_{j}}{\|\bm{f}^{g}_{i}\|\|\bm{f}^{g}_{j}\|}\cdot\left({\bm{f}^{g}_{i}}^{\top}\bm{f}^{g}_{j}{>}0\right),&\text{if }p^{g}_{j}{\in}\mbox{$k$-NN}(p^{g}_{i}),\\ 0,&\text{otherwise}.\end{cases}$
@@
^$\hat{\mathcal{F}}^{g}\leftarrow\text{ReLU}\left(W_{m}\left(\tilde{\mathcal{W}}^{g}\hat{\mathcal{F}}^{g}\right)+W_{s}\hat{\mathcal{F}}^{g}\right)$ (4)
@@ 3.2 Learning with local representation consensus
The main learning objective encourages the model to generate outputs that are close to the target responses in the training dataset. This objective is typically expressed as a per-token cross-entropy loss $\mathcal{L}_{pred}$ [9, 21] (see details in our Supp. Mat.). We found that adding a consensus term to $\mathcal{L}_{pred}$ that encourages local regularization of representations improves training stability. Firstly, points belonging to the same object (i.e.local neighborhoods) should share similar aggregated representations. Secondly, aggregated super-point representations should remain close to their original point representations to preserve context knowledge. Our novel consensus loss $\mathcal{L}_{\text{con}}$ is defined as
@@
$\mathcal{L}=\lambda{\mathcal{L}}_{con}+{\mathcal{L}}_{pred},$ (6)
@@ 4 Experiments
Implementation Details. As in [9, 12], we input 40,000 randomly sampled points from each 3D scene. We split the point cloud into $L=6$ partitions, and choose $\texttt{k}{=}4L{=}24$ neighbors to enhance global representations. We set $\lambda{=}\mu{=}0.1$ in the loss function. We use the pre-trained OPT-1.3B [62] language model, kept frozen and loaded in float16 precision. We use the AdamW [35] optimizer with a weight decay of 0.1, applying a cosine annealing scheduler that decays the learning rate from $10^{-4}$ to $10^{-6}$ over approximately 100,000 iterations. All tasks are trained with a total batch size of 16. Each training process is completed within two days using up to two NVIDIA H100 P0 (96GB) GPUs. For each evaluation, we fine-tune the model’s parameters on the respective task for about 30,000 iterations.
@@ 4.2 Ablation study
Local and global representations.
To investigate the impact of local and global representations, we design two variants of PerLA: one using only local representations and the other using only global representations. We also report an extended version of LL3DA, termed LL3DA, in which we increase the number of query tokens used to encode local partitions. Representations are first extracted from each partition, and FPS is applied to sample 1,024 points with their corresponding representations from the union of these partitions. These representations are then passed through a multimodal adapter to produce 32 global tokens. LL3DA generates an additional 32 global tokens from the point cloud of the entire scene, resulting in a total of 64 tokens, which are concatenated and fed into the LLM for response generation. (For details of LL3DA please refer to the Supp. Mat.). Tab. 3 shows that the combination of local and global representations yields a much better performance than the two independently used. Tab. 3 also shows that LL3DA does not reach PerLA performance, underlying the importance of information exchange between local and global representations.
Table 3: Ablation study of our local-to-global representation aggregation algorithm on ScanQA, ScanRefer[6] and Nr3D[1] benchmarks. LL3DA denotes an extended version of LL3DA with an increased number of query tokens.
@@
Localized cross-attention.
To evaluate the effectiveness of our novel localized cross-attention module, we tested max pooling and mean pooling as alternative methods. Tab. 4 shows that localized cross-attention significantly outperforms both mean pooling and max pooling across all metrics. These results indicate that localized cross-attention enhances ability of model to generate more informative scene representations, thereby improving performance on 3D question answering tasks. This improvement comes from the incorporation of positional information and semantic similarity in cross-attention, which helps to exclude neighboring points that do not belong to the same object.
@@
Loss function. To evaluate the effectiveness of our novel loss term $\mathcal{L}_{\text{con}}$, we train PerLA using different combinations of $\mathcal{L}_{\text{con}}$’s components (Eq. 5). Since our task is intended for downstream applications, each combination also includes the task-specific loss $\mathcal{L}_{\text{pre}}$. Tab. 5 shows that adding $\mathcal{L}_{\text{smt}}$ to $\mathcal{L}_{\text{pre}}$ improves all metrics (C, B4, M, and R). Incorporating the loss of regularization $\mathcal{L}_{\text{reg}}$ further enhances performance, particularly in the metrics C, M, and R. These results show the effectiveness of learning jointly with both semantic awareness and regularization losses.
@@
Number of partitions. To evaluate the effectiveness of different number of partitions, we test PerLA by splitting the point cloud into 4, 6, and 8 partitions. Tab. 6 shows that increasing the number of partitions generally improves performance, with 6 partitions yielding the best results. Although 8 partitions outperform 4, they show slightly lower performance than 6. This suggests that 6 partitions provide a better balance between granularity and effectiveness, capturing relevant scene details with lower computational complexity. Additional ablation studies, including an analysis of performance in different training strategies, can be found in Supp. Mat..

### v2

@@ Abstract
Enabling Large Language Models (LLMs) to understand the 3D physical world is an emerging yet challenging research direction. Current strategies for processing point clouds typically downsample the scene or divide it into smaller parts for separate analysis. However, both approaches risk losing key local details or global contextual information. In this paper, we introduce PerLA, a 3D language assistant designed to be more perceptive to both details and context, making visual representations more informative for the LLM. PerLA captures high-resolution (local) details in parallel from different point cloud areas and integrates them with (global) context obtained from a lower-resolution whole point cloud. We present a novel algorithm that preserves point cloud locality through the Hilbert curve and effectively aggregates local-to-global information via cross-attention and a graph neural network. Lastly, we introduce a novel loss for local representation consensus to promote training stability. PerLA outperforms state-of-the-art 3D language assistants, with gains of up to +1.34 CiDEr on ScanQA for question answering, and +4.22 on ScanRefer and +3.88 on Nr3D for dense captioning. Project page: https://gfmei.github.io/PerLA
@@ 2 Related work
3D language understanding involves understanding 3D scenes by describing or answering to scene-relevant questions in the format of natural language. Examples of typical downstream tasks include 3D Dense Captioning [8, 12, 56], and 3D Question Answering [2, 39, 65, 67]. 3D Dense Captioning consists of 3D localization of object instances, and textual descriptions of each object instance [12, 4, 8, 10]. 3D Question Answering requires the model equipped with a language decoder to answer questions regarding the visual context in the given 3D scene [2, 44]. Several works focus on addressing the problems of 3D-language pre-alignment [13, 26], or designing adapter layers [21, 9], or constructing 3D synthetic data [64]. In contrast, our work focuses on designing an encoding approach that can enhance 3D-language understanding by capturing fine-grained details.
3D language assistants (3DLAs) can be roughly categorized into two types: object-level 3DLAs [46, 53, 61, 23] and scene-level 3DLAs [21, 69, 59, 9]. Object-level 3DLAs learn from large 3D object datasets [15] to connect object-level 3D representation with language models. However, they underperform in compositional reasoning in complex 3D scenes with numerous objects. Scene-level 3DLAs, such as 3D-LLM [21], LLaVA-3D [69], Chat-3D [59], and LL3DA [9] enable scene understanding via the interaction with objects. Because 3D data is orders of magnitude less than 2D data, existing 3DLAs address such lack of data by leveraging pre-trained 2D Large Multimodal Models (LMMs) [21, 69], or data-efficient training recipes [59, 9]. 3D-LLM [21] leverages 2D pre-trained representations of rendered multi-view images to construct 3D representations and 2D VLMs as backbones. LLaVA-3D [69] adapts LLaVA [36] for 3D scene understanding by associating 2D patch representations with their positions in 3D space. Chat-Scene [22] improves the referencing and grounding capability and models scene representations as a sequence of object-level representations. LL3DA [9] extracts point-level representations from a downsampled 3D scene, and includes both interaction prompts and textual instructions to resemble human interactions with the 3D environment. Orthogonal to the efforts of aligning 3D-language with limited 3D data in existing 3DLAs, PerLA aims to improve the capability of 3DLAs in perceiving scene details.
@@ 3.1 Perceptive scene encoder
Global and partial scene encoding. Both full-scene point cloud and the partial-scene point clouds are encoded separately by the same pre-trained 3D scene encoder $\phi$. The 3D scene encoder $\phi$ downsamples the input point cloud into super-points using Farthest Point Sampling (FPS) [41] and produces a representation for each super-point. The representation of the full scene, i.e., the global representation, encode the overall scene context, while the representations of the partial scenes, i.e., the local representations, encode scene details. Let $\mathcal{P}^{g}=\{\bm{p}^{g}_{i}\in\mathbb{R}^{3}\}_{i=1}^{M}$ denote the downsampled full-scene point cloud composed of $M$ points, and $\mathcal{F}^{g}=\{\bm{f}^{g}_{i}\in\mathbb{R}^{d}\}_{i=1}^{M}$ the associated global representations. Let $\mathcal{P}^{l}=\{\bm{p}^{l}_{i}\in\mathbb{R}^{3}\}_{i=1}^{L\cdot M}$ denote the set of downsampled partial-scene point clouds, and $\mathcal{F}^{l}=\{\bm{f}^{l}_{i}\in\mathbb{R}^{d}\}^{L\cdot M}$ the associated local representations. Note that, we aggregate all the points of the downsampled partial-scene point clouds into the single set $\mathcal{P}^{l}$, and both the full-scene and the partial-scenes are downsampled to the same number of super-points $M$. The representations $\mathcal{F}^{l}$ provide a more detailed view of the scene, as they are derived from the same number of points downsampled from smaller, localized regions, resulting in higher resolution compared to the global versions. Yet, their semantic visibility is focused within each local part. In the following steps, we aggregate local and global information, to enrich the representations with details within the scene context.
@@
Hilbert-based nearest-neighbor search. In order to enhance global representations through local representations, we need to first find the correspondences among them, i.e., which super-points of partial-scene point clouds are spatially neighbors with super-points of the full-scene point cloud. This can be done via $k$-NN search. However, traditional $k$-NN searches can be computationally intensive on large-scale or high-resolution point clouds and cannot ensure that the identified neighbors maintain geometric consistency. To address these limitations, we approximate $k$-NN with an efficient neighbor mapping technique using Hilbert serialized point clouds, which improves query speed while reducing computational complexity. With serialized point clouds, we can leverage point indices to perform $k$-NN searches with $O(1)$ complexity [60].
@@
Specifically, we first serialize the union $\mathcal{P}^{g}\bigcup\mathcal{P}^{l}$ using the Hilbert curve ordering. We then apply geometric partitioning [27] on the original point clouds to generate geometric labels $\mathcal{Y}^{g}\bigcup\mathcal{Y}^{l}$. The center points of these labels are incorporated as high-order bits (by bit shifting) in the serialized index, ensuring that points with different labels occupy non-overlapping index ranges. Next, for each $\bm{p}^{g}_{i}\in\mathcal{P}^{g}$, we identify the k nearest super-points within $\mathcal{P}^{l}$ based on the serialized indexes. These geometric labels guarantee that $\bm{p}^{g}_{i}$ and its nearest local super-points all originate from the same instance. Let $\mathcal{P}_{\texttt{k}_{i}}^{l}$ denote the set of k nearest local super-points to the global super-point $\bm{p}^{g}_{i}$. Consequently, $\mathcal{F}_{\texttt{k}_{i}}^{l}$ are the representations of the super-points in $\mathcal{P}_{\texttt{k}_{i}}^{l}$.
@@
Localized cross-attention operates with a learning-based weighting mechanism that is a function of the representations and relative positions of global super-points and their associated local neighbors. We constrain cross-attention to local neighborhood regions because i) limiting the number of points reduces computational complexity, and ii) neighbors of a given point are likely to belong to the same object. Specifically, for each point $\bm{p}^{g}_{i}$ and its nearest neighbors $\mathcal{P}_{\texttt{k}_{i}}^{l}$, their relative position embeddings $\mathcal{R}_{i}$ are extracted as
$\vspace{-1mm}\mathcal{R}_{i}=\left\{\mathcal{R}_{ij}=\text{pos}\left(\left(\bm{p}^{g}_{i}-\bm{p}^{l}_{j}\right)/{\sigma}\right)\in\mathbb{R}^{d}\right\}_{j=1}^{K^{l}},$ (1)
@@
$\vspace{-2mm}\text{pos}(\bm{x})=\left[\sin\left(2\pi\bm{x}\cdot B\right);\cos\left(2\pi\bm{x}\cdot B\right)\right],$ (2)
@@
GCN message passing further refines ^$\hat{\mathcal{F}}^{g}$ by aggregating information from neighboring global super-points. Let $\mathcal{G}^{g}=\{\mathcal{P}^{g},\mathcal{W}^{g}\}$ be a $k$-NN graph, where $\mathcal{P}^{g}$ is the set of the vertices, i.e., global super-points, and $\mathcal{W}^{g}\in\mathbb{R}^{N\times N}$ is the adjacency matrix defining the edges. The construction of the adjacency matrix uses the similarity between super-point representations $\mathcal{F}^{G}$. Note that we use the (original) global representations to define the adjacency matrix because we empirically experienced more stability in training as opposed to using the updated global representation, likely due to the fact that they change value during training. The adjacency matrix $\mathcal{W}^{g}$ with elements $\mathcal{W}^{g}_{ij}$ is defined as
$\mathcal{W}^{g}_{ij}{=}\begin{cases}\frac{{\bm{f}^{g}_{i}}^{\top}\bm{f}^{g}_{j}}{\|\bm{f}^{g}_{i}\|\|\bm{f}^{g}_{j}\|}\cdot\left({\bm{f}^{g}_{i}}^{\top}\bm{f}^{g}_{j}{>}0\right),&\text{if }p^{g}_{j}{\in}\mbox{$k$-NN}(p^{g}_{i}),\\ 0,&\text{otherwise}.\end{cases}\vspace{-2mm}$
@@
^$\hat{\mathcal{F}}^{g}\leftarrow\text{ReLU}\left(W_{m}\left(\tilde{\mathcal{W}}^{g}\hat{\mathcal{F}}^{g}\right)+W_{s}\hat{\mathcal{F}}^{g}\right),\vspace{-2mm}$ (4)
@@ 3.2 Learning with local representation consensus
The main learning objective encourages the model to generate outputs that are close to the target responses in the training dataset. This objective is typically expressed as a per-token cross-entropy loss $\mathcal{L}_{pred}$ [9, 21] (see details in our Supp. Mat.). We found that adding a consensus term to $\mathcal{L}_{pred}$ that encourages local regularization of representations improves training stability. Firstly, points belonging to the same object (i.e., local neighborhoods) should share similar aggregated representations. Secondly, aggregated super-point representations should remain close to their original point representations to preserve context knowledge. Our novel consensus loss $\mathcal{L}_{\text{con}}$ is defined as
@@
$\mathcal{L}=\lambda{\mathcal{L}}_{con}+{\mathcal{L}}_{pred},\vspace{-2mm}$ (6)
@@ 4 Experiments
Implementation details. As in [9, 12], we input 40,000 randomly sampled points from each 3D scene. We divide the point cloud into partitions $L{=}6$ and choose $\texttt{k}{=}4L{=}24$ neighbors to enhance global representations. We set $\lambda{=}\mu{=}0.1$ in the loss. We use the pre-trained OPT-1.3B [66] language model, kept frozen and loaded in float16 precision. We use the AdamW [38] optimizer with a weight decay of 0.1, applying a cosine annealing scheduler that decays the learning rate from $10^{-4}$ to $10^{-6}$ over approximately 100,000 iterations. All tasks are trained with a total batch size of 16. Each training process is completed within two days using up to two NVIDIA H100 P0 (96GB) GPUs. For each evaluation, we fine-tune the model’s parameters on the respective task for about 30,000 iterations.
@@ 4.2 Ablation study
Table 3: Ablation study of our local-to-global, GCN representation aggregation algorithms on ScanQA, ScanRefer[6] and Nr3D[1] benchmarks. LL3DA denotes an extended version of LL3DA with an increased number of query tokens.
@@
Local, global, and GCN representations. To investigate the impact of local and global representations, we design two variants of PerLA: one using only local representations and the other using only global representations. We also report an extended version of LL3DA, termed LL3DA, in which we increase the number of query tokens used to encode local partitions. Representations are first extracted from each partition, and FPS is applied to sample 1,024 points with their corresponding representations from the union of these partitions. These representations are then passed through a multimodal adapter to produce 32 global tokens. LL3DA generates an additional 32 global tokens from the point cloud of the entire scene, resulting in a total of 64 tokens, which are concatenated and fed into the LLM for response generation. (For details of LL3DA please refer to the Supp. Mat.). Tab. 3 shows that the combination of local and global representations yields a much better performance than the two independently used. Tab. 3 also shows that LL3DA does not reach PerLA performance, underlying the importance of information exchange between local and global representations. To evaluate the effectiveness of GCN, we also report the results without using GCN (w/o GCN). Compared to our PerLA, which incorporates GCN, it consistently improves performance across the three datasets, confirming its effectiveness.
Localized cross-attention. To evaluate the effectiveness of our novel localized cross-attention module, we tested max pooling and mean pooling as alternative methods. Tab. 4 shows that localized cross-attention significantly outperforms both mean pooling and max pooling across all metrics. These results indicate that localized cross-attention enhances ability of model to generate more informative scene representations, thereby improving performance on 3D question answering tasks. This improvement comes from the incorporation of positional information and semantic similarity in cross-attention, which helps to exclude neighboring points that do not belong to the same object.
@@
Loss function. To evaluate the effectiveness of our novel loss term $\mathcal{L}_{\text{con}}$, we train PerLA using different combinations of $\mathcal{L}_{\text{con}}$’s components (Eq. 5). Since our task is intended for downstream applications, each combination also includes the task-specific loss $\mathcal{L}_{\text{pre}}$. Tab. 5 shows that adding $\mathcal{L}_{\text{smt}}$ to $\mathcal{L}_{\text{pre}}$ improves all metrics (C, B4, M, and R). Incorporating the loss of regularization $\mathcal{L}_{\text{reg}}$ further enhances performance, particularly in metrics C, M, and R. These results show the effectiveness of learning jointly with both semantic awareness and regularization losses.
@@
Number of partitions. To evaluate the effectiveness of different number of partitions, we test PerLA by splitting the point cloud into 4, 6, and 8 partitions. Tab. 6 shows that increasing the number of partitions generally improves performance, with 6 partitions yielding the best results. Although 8 partitions outperform 4, they show slightly lower performance than 6. This suggests that 6 partitions provide a better balance between granularity and effectiveness, capturing relevant scene details with lower computational complexity. More ablation studies, including an analysis of performance in different training strategies, refer to Supp. Mat..